Reports until 16:38, Tuesday 28 January 2025
GQuEST Output Filter Cavities
Torrey Cullen - posted 16:38, Tuesday 28 January 2025 - last comment - 15:10, Wednesday 29 January 2025(12099)
Lab Work 1/28 - Continued OFC3 Work

[Jeff,Torrey]

Assorted notes:

-I completed the assembly of the piezo mirror. This required having daniel machine a custom spacer for the assembly. With the spacer, the piezo alignment tool works. Note that this new piezo has a larger OD that is large enough to prohibit the use of the nylon tipped set screws for alignment. Because of this the spacer and piezo may not be concentrically aligned to the mirror.

-Installed a 2 meter 780 PM fiber at the OFC2/OFC3 sled fiber port hub. This allows a connection from the third AOM path to OFC3 775 input.

-We are going to start mode matching into cavity 3. The USB extension from the NUC computers does not recognize the scanning slit beam profiler. This is probably due to the USB extension being over 100 ft and needing external power. We are going to use the new lab laptop for now instead.

-The new laptop is unusable for the scanning slit software. For some reason it is constantly crashing and freezing.

-After switching to Torrey's laptop, we collected data for the beam size between the launcher and the third filter cavity. We noticed amplitude fluctuations on the beam, but they reduced when we raised the power on this path (by turning the amplifier current to 1A. This should correspond to ~1.5W total output on the distribution center).

-We spent some time profiling and finding a mode matching solution. The waist for this light needs to be .26 meters after the collimator for this configuration which is making it difficult to mode match. We are going to try moving the collimator back tomorrow and reprofiling.

-We turned the amplifier back down until tomorrow.

Comments related to this report
Torrey Cullen - 12:05, Wednesday 29 January 2025 (12100)

After playing around in JAMMT I found that moving the collimator back would result in easy mode matching solutions. I moved the collimator back 12 inches and reprofiled. Now, according to finesse and JAMMT, a robust solution is f=2000 mm @ z=0.066, f=2500 mm @ z=0.354, (v=100 %). We have these lens. Will implement after lunch.

Torrey Cullen - 15:10, Wednesday 29 January 2025 (12102)

I put the two lens in the positions according to the above comment. I then profiled the beam after the second lens. This gives the following data below:

                                         Horizontal    Vertical
Beam waist radius[um]    402.531593    423.098793
Waist position [m]            0.148341    0.146998
Rayleigh range [inches]   25.859159    28.569197


Radius Average               412.815193
Waist Position Average    0.147669

According to finesse, and accounting for the extra two 1/8 spacers, the beam this cavity accepts is 411 um @ .106 meters after the lens. The measured quanity is 412 um .147 meters after the lens. This should yield a mode mismatch well below 1%. I will close this work for now unless we see large laguerre gauss modes in the cavity scans in the future.

Images attached to this comment